OpenStudy (anonymous):

square root of 72 over square root of 108

OpenStudy (diyadiya):

\[\sqrt{\frac{2}{3}}\]

OpenStudy (anonymous):

Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 40 ÷ 5(3 - 1) - 4 Line 2 40 ÷ 5(2) - 4 Line 3 40 ÷ 10 - 4 Line 4 4 - 4 Line 5 0

jhonyy9 (jhonyy9):

sqrt(72/108)=sqrt(8/12)=sqrt(2/3)

OpenStudy (anonymous):

you cant have a square root on the bottom

jhonyy9 (jhonyy9):

than is sqrt6/3